OpenStudy (anonymous):

Calculate dy/dx. You need not expand your answer. y = (x + 4)(x^2 − 4)

OpenStudy (mimi_x3):

Use the product rule. u = (x+4) u'= 1 v = x^2-4 v' = 2x Sub it into the product rule and solve.